OpenStudy (anonymous):

Find the derivative of [cos(x)^2][sin(x)] and it has to equal cos(x)[1-3sin(x)^2]. Also you have to use the identity sin(x)^2+cos(x)^2=1. So far I have: Using product rule- u'(x)v(x)+v'(x)u(x) =[(cos'(x)^2)*sin(x)] + [sin'(x)*cos(x)^2] using chain rule =[2cos(x)*cos'(x)*sin(x)] + [cos(x)*cos(x)^2] simplifying =2cos(x)[sin(x)^2]+cos(x)^3 And after this I don't know what to do, but I may have messed up on a previous step or have overlooked an identity I could have used.

OpenStudy (anonymous):

use the identity you posted sin(x)^2+cos(x)^2=1 and rearrange for cos(x)^2 cos(x)^2 = 1-sin(x)^2 plug it in: [sin(x)]1-sin(x)^2 = sin(x) - sin(x)^3 take the derivative: cos(x) -3sin(x)^2 <---- answer

OpenStudy (anonymous):

I don't understand how you got sin(x)^3. Also I kind of learned it that you set it up in product/quotient/etc before you do any identities. Finally, when you took the derivative, it needs to be [1-3sin(x)^2] and so I think a one got dropped off at some point.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

Do you know the distributive property? @mmorrison52